OpenStudy (nikolas):

(imaginary number) does (-i)^2 = -1? That is what I have in my notes but I think it should be +1, by the following reasoning: (-i)^2=(-sqrt{-1})^2=-(-1)=+1 Am I correct and just made a mistake in my notes, or have I completely misunderstood something? Hope someone with a little more experience can help clarify this. Thanks!

hartnn (hartnn):

\((-i)^2= (-1)^2i^2 = 1 (-1)=-1\)

hartnn (hartnn):

from your way, (-sqrt{-1})^2 = (-1)^2 * (sqrt{-1})^2 = 1 * (-1) = -1

OpenStudy (nikolas):

Ah, can't believe I missed separating the (-1). Thanks for your help!

hartnn (hartnn):

welcome ^_^